Find Mx and My for the first quadrant area bounded by x^2=y^2=a^2 and the coordinate axes


1
Expert's answer
2021-08-25T12:11:18-0400

The set for which you need to find "M_x" , "M_y" is "\{(x;y)| 0\leq x\leq a, 0\leq y\leq a\}" .

"M_x=p\int\limits_a^b\frac{1}{2}[f^2(x)-g^2(x)]dx,""M_y=p\int\limits_a^bx[f(x)-g(x)]dx" . In this case "f(x)=a" and "g(x)=0" . So "M_x=p\int\limits_0^a \frac{1}{2}a^2dx=\frac{a^2p}{2}(a-0)=\frac{1}{2}pa^3" and "M_y=p\int\limits_0^a xadx=\frac{1}{2}pa^3"